# GrokAlign: Geometric Characterisation and Acceleration of Grokking

### v1

@@ 4 Experiments
Figure 3: Theorem 9 holds in practice: we observe that a changing inner product indeed corresponds to a feature learning regime. Here we train a two-layer scalar-output ReLU network using the binary-cross-entropy loss function to distinguish between the zero and one class of the MNIST dataset [1]. We train the model using full-batch gradient descent for $4000$ steps at a learning rate of $0.01$. At the beginning of training we fix a point from the training set and compute the average value of the neural tangent kernel between itself and the other points from the training set, left. We then compute the centroid of the point using Theorem 4 to then obtain the inner product, centre, and its alignment, right. In the right plot we record the norm of the centroid and encode it in the colours of the markers. In the left and centre plots, we identify the feature learning regime using a green shaded area, determined by the neural tangent kernel changing value.
@@ 4.1 Centroid Alignment Identifies the Feature Learning Regime
Figure 4: Centroid alignment of a point from the training set identifies the generalisation and robustness of a deep network. We study a two-layer network of width $2048$ learning the XOR classification task described in Section 4.2. In the top row we train the network using full-batch gradient descent under the mean-squared error loss function with a learning rate of $0.1$, weight-decay of $0.1$ and for $1000$ steps. We monitor both the alignment of a training point to its centroid as well as the robustness of the network. To measure robustness, we take the entries from the test set and apply Gaussian perturbations of varying standard deviations to the last $39998$ components. In the bottom row we additionally apply GrokAlign with $\lambda_{\text{Jac}}$ equal to $1.0$. In the left plots we are illustrating the centroid norms with the colours of the markers.
@@ 4.2 Centroid Alignment Identifies Delayed Robustness
We adopt a set-up similar to that of Xu et al. [36] entailing a scalar-output two-layer fully connected network grokking on XOR cluster data. Note that this network trivially has rank-one Jacobians at every point in the input space. The XOR cluster data contains $40000$-dimensional vectors of the form $\mathbf{x}=\left(x_{1},x_{2},\tilde{\mathbf{x}}^{\top}\right)^{\top}\in\mathbb{R}^{40000}$, where $x_{1},x_{2}\in\left\{\pm 1\right\}$ and $\tilde{\mathbf{x}}\in\mathbb{R}^{39998}$. The $400$ samples used to train the network are constructed by sampling entries $x_{1}$, $x_{2}$ uniformly from {±$\left\{\pm 1\right\}$ and entries of $\tilde{\mathbf{x}}$ uniformly from {±$\left\{\pm\epsilon\right\}$, here we take $\epsilon=0.05$. The corresponding label of such a sample is $x_{1}x_{2}\in\{\pm 1\}$. A similar sample is generated as a test set. Therefore, by construction our training data only has signal in the first two components, whereas all the other components contain noise. Hence, generalisation would require recognising the pattern of how the first two components lead to the corresponding label, whilst robustness would require the network to not condition its pattern recognition on the last $39998$ components.
@@ 4.3 GrokAlign’s Influence on Grokking
Figure 5: GrokAlign capitalises on the low rank implicit bias of deep network training to induce robustness by aligning the Jacobians of the deep network. Here we train a convolutional neural network with six convolutional layers and two linear layers, with no bias terms, on a $1024$ subset of the CIFAR10 dataset [37] under the mean-squared error loss function. We use the AdamW optimizer [29] with a learning of $0.001$ and a batch size of $256$ to train the network across $36000$ steps. In one instance we apply weight-decay at $0.001$, and in another instance we apply GrokAlign with $\lambda_{\text{Jac}}$ equal to $0.001$. In the left plot we compute the average explained variance of the first principal component of the Jacobians as in Figure 2, and in the centre plot we record the average centroid alignment on the training set. In the right column we record the test accuracy of the model and the accuracy of the model when $\ell_{\infty}$ perturbations of amplitude $\frac{4}{255}$ are applied to the test set using Autoattack [38].
@@
Figure 7: By maintaining the Frobenius norm of the Jacobians at the training data relatively high we can keep the norms of the centroids relatively high which prevents grokking. We take the MNIST grokking set up of Liu et al. [6]. In the minimising case we impose GrokAlign with $\lambda_{\text{Jac}}$ equal to $0.001$ during training to minimise the Frobenius norm of the Jacobians at the training data. In the constrained case we apply GrokAlign to maintain the Frobenius norm of the Jacobians computed at the training data at a relatively high level. More specifically, using a regularisation coefficient of $0.001$, we append the difference between five and the average Frobenius norm of the training data to the loss function. In the left plot we visualise the train and test accuracy with solid and dashed lines respectively. In the centre plot we visualise the average norm of the centroids computed at the training data. In the right plot we visualise the average centroid alignment.
@@
From Table 1, we observe that GrokAlign is extremely effective at inducing the grokked state of the deep network in this setting; it arrives at the grokked state in $7.56$ times fewer steps and $6.31$ times faster than the baseline in the case of the cross-entropy loss function. In contrast, Grokfast provides a relatively lower improvement in the mean-squared error case and is ineffective in the cross-entropy case. Adversarial training does not improve the rate of grokking over the baseline. In particular, adversarial training does not improve the robustness of the model by way of aligning the Jacobian, unlike GrokAlign (see Appendix C).
Table 1: GrokAlign significantly speeds up the rate of grokking. In the top table we assume the MNIST set up of Liu et al. [6] with the cross-entropy loss function, and in the bottom table we assume it with mean-squared error loss function as the baselines. We repeat the training across ten different random initialisations, where for the cross-entropy loss function we apply GrokAlign with $\lambda_{\text{Jac}}$ equal to $0.001$ and with $\lambda_{\text{Jac}}$ equal to $0.0001$ for the mean-squared error models. For each run we measure the number of steps and the absolute time, in seconds, taken for the networks to reach $85$% test accuracy. In the case of the mean-squared error loss function, we additionally measure the time, in seconds, for the models to go from $20$% to $85$% test accuracy. We provide the average acceleration (or deceleration) of each method compared to the baseline along with the corresponding standard deviation.
@@
Grokfast $\uparrow 1.01\times(\pm 0.04)$ $\uparrow 1.03\times(\pm 0.04)$
Adversarial Training $\downarrow 1.32\times(\pm 0.18)$ $\uparrow 1.92\times(\pm 0.29)$
@@
Grokfast $\downarrow 1.08\times(\pm 0.17)$ $\downarrow 1.05\times(\pm 0.23)$ $\downarrow 1.05\times(\pm 0.27)$
Adversarial Training $\downarrow 1.23\times(\pm 0.26)$ $\uparrow 1.92\times(\pm 0.29)$ $\uparrow 2.02\times(\pm 0.37)$
@@
Figure 8: GrokAlign biases a single layer transformer model to learn a classification style solution. Here we obtain the centroid alignment statistics for a single layer transformer trained on modular addition [4]. We use the same training pipeline as in Nanda et al. [15], with the added utilisation of GrokAlign with $\lambda_{\text{Jac}}$ equal to $0.001$. In the left plot we record the accuracies of the transformer model on the train and test set. In the centre plot we record the centroid alignment of the transformer model from the continuous embeddings of the discrete input tokens to the output of the model; the shaded region illustrates the maximum and minimum alignment observed with the solid line corresponding to the mean. In the right plot we record the Gini coefficients of the embedding and unembedding matrices of the model as proposed in Nanda et al. [15].
@@ Appendix C Alignment Induced by Adversarial Training
Regularisation Test Accuracy Jacobian Row Alignment (max/min) Baseline $88.9$% $-0.254$/$0.283$ GrokAlign $\mathbf{91.8}$$\mathbf{-0.509}$/$\mathbf{0.478}$ Adversarial Training $88.0$% $-0.253$/$0.274$
Regularisation Test Accuracy Jacobian Row Alignment (max/min) Baseline $86.8$% $-0.170$/$0.272$ GrokAlign $\mathbf{88.2}$$\mathbf{-0.638}$/$\mathbf{0.709}$ Adversarial Training $87.8$% $-0.263$/$0.376$

### v2

@@ 4 Experiments
Figure 3: Theorem 9 holds in practice: we observe that a changing inner product indeed corresponds to a feature learning regime. Here we train a two-layer scalar-output ReLU network using the binary-cross-entropy loss function to distinguish between the zero and one class of the MNIST dataset [1]. We train the model using full-batch gradient descent for 400040004000 steps at a learning rate of 0.010.010.01. At the beginning of training we fix a point from the training set and compute the average value of the neural tangent kernel between itself and the other points from the training set, left. We then compute the centroid of the point using Theorem 4 to then obtain the inner product, centre, and its alignment, right. In the right plot we record the norm of the centroid and encode it in the colours of the markers. In the left and centre plots, we identify the feature learning regime using a green shaded area, determined by the neural tangent kernel changing value.
@@ 4.1 Centroid Alignment Identifies the Feature Learning Regime
Figure 4: Centroid alignment of a point from the training set identifies the generalisation and robustness of a deep network. We study a two-layer network of width 204820482048 learning the XOR classification task described in Section 4.2. In the top row we train the network using full-batch gradient descent under the mean-squared error loss function with a learning rate of 0.10.10.1, weight-decay of 0.10.10.1 and for 100010001000 steps. We monitor both the alignment of a training point to its centroid as well as the robustness of the network. To measure robustness, we take the entries from the test set and apply Gaussian perturbations of varying standard deviations to the last 399983999839998 components. In the bottom row we additionally apply GrokAlign with $\lambda_{\text{Jac}}$ equal to 1.01.01.0. In the left plots we are illustrating the centroid norms with the colours of the markers.
@@ 4.2 Centroid Alignment Identifies Delayed Robustness
[…]
@@ 4.3 GrokAlign’s Influence on Grokking
Figure 5: GrokAlign capitalises on the low rank implicit bias of deep network training to induce robustness by aligning the Jacobians of the deep network. Here we train a convolutional neural network with six convolutional layers and two linear layers, with no bias terms, on a 102410241024 subset of the CIFAR10 dataset [37] under the mean-squared error loss function. We use the AdamW optimizer [29] with a learning of 0.0010.0010.001 and a batch size of 256256256 to train the network across 360003600036000 steps. In one instance we apply weight-decay at 0.0010.0010.001, and in another instance we apply GrokAlign with $\lambda_{\text{Jac}}$ equal to 0.0010.0010.001. In the left plot we compute the average explained variance of the first principal component of the Jacobians as in Figure 2, and in the centre plot we record the average centroid alignment on the training set. In the right column we record the test accuracy of the model and the accuracy of the model when $\ell_{\infty}$ perturbations of amplitude $\frac{4}{255}$ are applied to the test set using Autoattack [38].
@@
Figure 7: By maintaining the Frobenius norm of the Jacobians at the training data relatively high we can keep the norms of the centroids relatively high which prevents grokking. We take the MNIST grokking set up of Liu et al. [6]. In the minimising case we impose GrokAlign with $\lambda_{\text{Jac}}$ equal to 0.0010.0010.001 during training to minimise the Frobenius norm of the Jacobians at the training data. In the constrained case we apply GrokAlign to maintain the Frobenius norm of the Jacobians computed at the training data at a relatively high level. More specifically, using a regularisation coefficient of 0.0010.0010.001, we append the difference between five and the average Frobenius norm of the training data to the loss function. In the left plot we visualise the train and test accuracy with solid and dashed lines respectively. In the centre plot we visualise the average norm of the centroids computed at the training data. In the right plot we visualise the average centroid alignment.
@@
From Table 1, we observe that GrokAlign is extremely effective at inducing the grokked state of the deep network in this setting; it arrives at the grokked state in 7.567.567.56 times fewer steps and 6.316.316.31 times faster than the baseline in the case of the cross-entropy loss function. In contrast, Grokfast provides a relatively lower improvement in the mean-squared error case and is ineffective in the cross-entropy case. Adversarial training does not improve the rate of grokking over the baseline. In particular, adversarial training does not improve the robustness of the model by way of aligning the Jacobian, unlike GrokAlign (see Appendix C).
Table 1: GrokAlign significantly speeds up the rate of grokking. In the top table we assume the MNIST set up of Liu et al. [6] with the cross-entropy loss function, and in the bottom table we assume it with mean-squared error loss function as the baselines. We repeat the training across ten different random initialisations, where for the cross-entropy loss function we apply GrokAlign with $\lambda_{\text{Jac}}$ equal to 0.0010.0010.001 and with $\lambda_{\text{Jac}}$ equal to 0.00010.00010.0001 for the mean-squared error models. For each run we measure the number of steps and the absolute time, in seconds, taken for the networks to reach 858585% test accuracy. In the case of the mean-squared error loss function, we additionally measure the time, in seconds, for the models to go from 202020% to 858585% test accuracy. We provide the average acceleration (or deceleration) of each method compared to the baseline along with the corresponding standard deviation.
@@
Grokfast 1.01×(±0.04$\uparrow$ 1.01$\times$($\pm$ 0.04)↑ 1.01 × ( ± 0.04 ) 1.03×(±0.04$\uparrow$ 1.03$\times$($\pm$ 0.04)↑ 1.03 × ( ± 0.04 )
Adversarial Training 1.32×(±0.18$\downarrow$ 1.32$\times$($\pm$ 0.18)↓ 1.32 × ( ± 0.18 ) 1.92×(±0.29$\uparrow$ 1.92$\times$($\pm$ 0.29)↑ 1.92 × ( ± 0.29 )
@@
Grokfast 1.08×(±0.17$\downarrow$ 1.08$\times$($\pm$ 0.17)↓ 1.08 × ( ± 0.17 ) 1.05×(±0.23$\downarrow$ 1.05$\times$($\pm$ 0.23)↓ 1.05 × ( ± 0.23 ) 1.05×(±0.27$\downarrow$ 1.05$\times$($\pm$ 0.27)↓ 1.05 × ( ± 0.27 )
Adversarial Training 1.23×(±0.26$\downarrow$ 1.23$\times$($\pm$ 0.26)↓ 1.23 × ( ± 0.26 ) 1.92×(±0.29$\uparrow$ 1.92$\times$($\pm$ 0.29)↑ 1.92 × ( ± 0.29 ) 2.02×(±0.37$\uparrow$ 2.02$\times$($\pm$ 0.37)↑ 2.02 × ( ± 0.37 )
@@
Figure 8: GrokAlign biases a single layer transformer model to learn a classification style solution. Here we obtain the centroid alignment statistics for a single layer transformer trained on modular addition [4]. We use the same training pipeline as in Nanda et al. [15], with the added utilisation of GrokAlign with $\lambda_{\text{Jac}}$ equal to 0.0010.0010.001. In the left plot we record the accuracies of the transformer model on the train and test set. In the centre plot we record the centroid alignment of the transformer model from the continuous embeddings of the discrete input tokens to the output of the model; the shaded region illustrates the maximum and minimum alignment observed with the solid line corresponding to the mean. In the right plot we record the Gini coefficients of the embedding and unembedding matrices of the model as proposed in Nanda et al. [15].
@@ 5 Discussion
Acknowledgments and Disclosure of Funding
This work was supported by ONR grant N00014-23-1-2714, ONR MURI N00014-20-1-2787, DOE grant DE-SC0020345, and DOI grant 140D0423C0076.
@@ Appendix C Alignment Induced by Adversarial Training
Regularisation Test Accuracy Jacobian Row Alignment (max/min) Baseline 88.988.988.9% 0.254-0.254- 0.254/0.2830.2830.283 GrokAlign $\mathbf{91.8}$$\mathbf{-0.509}$/$\mathbf{0.478}$ Adversarial Training 88.088.088.0% 0.253-0.253- 0.253/0.2740.2740.274
Regularisation Test Accuracy Jacobian Row Alignment (max/min) Baseline 86.886.886.8% 0.170-0.170- 0.170/0.2720.2720.272 GrokAlign $\mathbf{88.2}$$\mathbf{-0.638}$/$\mathbf{0.709}$ Adversarial Training 87.887.887.8% 0.263-0.263- 0.263/0.3760.3760.376